# Oriented matroids and type $\mathbb{A}$ cluster categories

### v2

@@ Abstract.
I would like to thank Rudradip Biswas for the opportunity to speak about a preliminary version of this work at the “New Directions in Group Theory and Triangulated Categories” seminar, and Mikhail Gorsky for the reference [8]. I am currently supported by EPSRC grant EP/W001780/1.
@@ 1. Introduction
From the very beginning, there has been an intimate relationship between cluster algebras and the combinatorics of triangulations. The most fundamental example of this is the $\mathbb{A}_{n}$ cluster algebra [10], whose cluster variables are in bijection with arcs in a convex $(n+3)$-gon, with clusters in bijection with triangulations, and mutation corresponding to flipping an arc in a quadrilateral. This was extended to cluster algebras from triangulations of surfaces in [9].
Higher-even-dimensional triangulations were connected with cluster algebras in [16], and higher odd dimensions in [22]. In particular, in [22] it was shown that equivalence classes of maximal green sequences of a linearly oriented type $\mathbb{A}$ quiver were in bijection with triangulations of a three-dimensional cyclic polytope. Maximal green sequences are finite sequences of mutations in a cluster algebra, first introduced into the mathematics literature by Keller [13], and into the physics literature by Cecotti, Cordova, and Vafa [6, Section 4.3]. In fact, these physicists already made a connection between maximal green sequences and three-dimensional triangulations [6, Section 5]. This appearance of maximal green sequences in physics is related to their role in quantum dilogarithm identities [20] and the Kontsevich–Soibelman wall-crossing formula from Donaldson–Thomas theory [14].
@@
Given the bijection from [22] between triangulations of three-dimensional cyclic polytopes and equivalence classes of maximal green sequences of linearly-oriented $\mathbb{A}_{n}$, the following natural question then arises. Given an arbitrary orientation of the type $\mathbb{A}$ Dynkin diagram, is there a three-dimensional polytope whose triangulations are in bijection with its maximal green sequences, up to equivalence? Following [22], this bijection should be such that the simplices in the polytope are in bijection with the mutations in the maximal green sequence, so that the number of simplices of the triangulation equals the length of the maximal green sequence.
One can deduce what properties such a three-dimensional polytope should have. An exchange pair in a maximal green sequence corresponds to flipping an arc in a quadrilateral. As in [22], one should lift this quadrilateral to a $3$-simplex such that the green mutation goes from the lower side of the simplex to the upper side. This amounts to choosing an orientation for each $3$-simplex in our three-dimensional polytope, and this determines the corresponding oriented matroid.
In order to determine the orientation of each $3$-simplex, we need a criterion which tells us, given an initial cluster, which mutations are green. Such a criterion is given by [8] in terms of the cluster category of [4], where a choice of initial cluster is given by a choice of cluster-tilting object $\mathsf{T}$. In particular, different orientations of $\mathbb{A}_{n}$ can be achieved by different choices of $\mathsf{T}$. Of course, it is not obvious that orienting the $3$-simplices according to which exchange pairs are made green by the cluster-tilting object $\mathsf{T}$ gives a well-defined oriented matroid $\mathcal{M}_{\mathsf{T}}$: certain axioms need to be satisfied. The first main theorem of this paper is that these axioms do in fact hold. We moreover conjecture that $\mathcal{M}_{\mathsf{T}}$ is indeed the oriented matroid of some polytope: that $\mathcal{M}_{\mathsf{T}}$ is “realisable”.
@@
The oriented matroids $\mathcal{M}_{\mathsf{T}}$ are orientations of the uniform matroid: all possible $3$-simplices are bases of the matroid. Orientations of the uniform matroid have been studied recently in the context of their associated “chirotropical Grassmannians”, which were introduced in [5] in the context of scattering amplitudes, and subsequently studied in [1]. While uniform matroids are always realisable [2, Section 8.3], uniform oriented matroids are not always realisable [2, Proposition 8.3.2].
@@ 2.1. Oriented matroids
A matroid is a structure abstracting the linear dependence relations that hold among a finite set of vectors. An oriented matroid is a matroid with some extra structure — an “orientation” — which also keeps track of the signs of the coefficients in these linear relations. We will follow the standard reference for oriented matroids [2].
@@
    $\chi(x_{\pi(1)},x_{\pi(2)},\dots,x_{\pi(r)})=\mathsf{sign}(\pi)(x_{1},x_{2},\dots,x_{r})$
@@
The chirotope $\chi$ is realisable if there exists a map $\rho\colon E\to\mathbb{R}^{r}$ such that
@@
Aside from the definition of matroids using bases, one can also define them using “circuits”. In terms of linear algebra, circuits are minimal linearly dependent sets. That is, a circuit is a linearly dependent set whose subsets are all linearly independent. Oriented matroids have signed circuits, which are signed sets.
Definition 2.2 ([2, p.102]).
A signed set $X$ is a set $\underline{X}$ together with a partition $(X^{+},X^{-})$ of $\underline{X}$ into two distinguished subsets. Here $\underline{X}$ is called the support of $X$. Given a set $E$, a signed subset of $E$ is a signed set $X$ with $\underline{X}\subseteq E$.
One can specify axioms for a set of signed subsets of the ground set $E$ to be the set of signed circuits of an oriented matroid [2, Definition 3.2.1]. Note that we do not identify an oriented matroid with its chirotope or with its set of signed circuits. We think of oriented matroids as objects in the style of object-oriented programming: an oriented matroid has a chirotope and a set of signed circuits as attributes; it is determined by specifying either of these pieces of data, but it is not identified with either of them. We say that an oriented matroid is realisable if its chirotope is realisable.
@@
Let $\chi\colon E^{r}\to\{-1,0,+1\}$ be the chirotope of $\mathcal{M}$. If $\mathcal{M}(V)$ also has rank $r$, then its chirotope is $\chi|_{V}$.
For every oriented matroid $\mathcal{M}$ on a set $E$, there is a dual oriented matroid $\mathcal{M}^{\ast}$ on $E$ [2, Proposition 3.4.1]. The circuits of $\mathcal{M}^{\ast}$ are known as the cocircuits of $\mathcal{M}$. The set of cocircuits of $\mathcal{M}$ is denoted $\mathcal{C}^{\ast}$. If $\mathcal{M}$ is realisable, then cocircuits of $\mathcal{M}$ correspond to hyperplanes, with one half of the cocircuit lying on one side of the hyperplane, the other half lying on the other [2, p.116], and the complement of the cocircuit lying in the hyperplane.
Definition 2.4 ([2, p.377]).
@@
Let $\mathcal{M}$ be an acyclic oriented matroid on a set $E$. A hyperplane of $\mathcal{M}$ is a subset $H\subseteq E$ such that $E\setminus H$ is a cocircuit of $\mathcal{M}$. An open halfspace of $\mathcal{M}$ is $Y^{+}$ for $Y$ a cocircuit of $\mathcal{M}$. A facet of $\mathcal{M}$ is a hyperplane $H$ such that $E\setminus H$ is an open halfspace. A face of $\mathcal{M}$ is an intersection of facets of $\mathcal{M}$.
@@
    If $\sigma,\tau\in\Delta$, then $\sigma\cap\tau$ is a common face of the restrictions $\mathcal{M}(\sigma)$ and $\mathcal{M}(\tau)$.
  2. (2)
    If $\sigma\in\Delta$, then each facet of $\mathcal{M}(\sigma)$ is either a facet of $\mathcal{M}$ or is contained in precisely two cells of $\Delta$.
@@ 2.2. Cluster categories of type $\mathbb{A}$
Cluster categories were introduced in the paper [4], which we refer to for full background. Given an acyclic quiver $Q$, its path algebra is denoted $KQ$, where $K$ is a field. We write $\mathscr{D}^{b}(KQ)$ for the bounded derived category of right $KQ$-modules which are finite-dimensional over $K$. The cluster category $\mathscr{C}_{Q}$ of $Q$ is the orbit category $\mathscr{D}^{b}(KQ)/\nu\Sigma^{-2}$, where $\nu:=D(KQ)\otimes^{\mathbf{L}}_{KQ}-\colon\mathscr{D}^{b}(KQ)\to\mathscr{D}^{b}(KQ)$ is the Nakayama functor and $\Sigma$ is the suspension of $\mathscr{D}^{b}(KQ)$. Here $D(-)=\operatorname{Hom}_{K}(-,K)$ is the standard duality.
The linearly oriented $\mathbb{A}_{n}$ quiver is $1\to 2\to\dots\to n$ and we denote its cluster category by $\mathscr{C}_{n}$. In fact, $\mathscr{C}_{n}$ is independent of the choice of orientation of the quiver. The cluster category $\mathscr{C}_{n}$ is a $K$-linear triangulated category with finitely many indecomposable objects, and one can describe its morphisms and suspension functor combinatorially. Indeed, up to isomorphism, the indecomposable objects of $\mathscr{C}_{n}$ may be labelled $\mathsf{O}_{B}$ for $B=\{b_{1},b_{2}\}\in\binom{[n+3]}{2}$ with $b_{1}\neq b_{2}\pm 1$. Here, when we write $b_{2}\pm 1$, we are using modular arithmetic with representatives in $[n+3]$, rather than $\{0,1,\dots,n+2\}$, as is more usual; we will continue do this in this paper. Hence, indecomposable objects in $\mathscr{C}_{n}$ are in bijection with arcs in a convex $(n+3)$-gon.
We call a permutation in the subgroup of the symmetric group $\mathfrak{S}_{m}$ generated by the permutation $x\mapsto x+1$ a cyclic permutation. By rotation, we mean applying a cyclic permutation. Given $\{a_{1},a_{2},\dots,a_{l}\}\subseteq[m]$ with $l\geqslant 3$, we write $a_{1}\prec a_{2}\prec\dots\prec a_{l}$ if there is a cyclic permutation $\pi\in\mathfrak{S}_{m}$ such that $\pi(a_{1})<\pi(a_{2})<\dots<\pi(a_{l})$ in the usual ordering on $[m]$. We call the ordering $a_{1}\prec a_{2}\prec\dots\prec a_{l}$ a cyclic ordering. We then have that $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\neq 0$ if and only if $b_{1}-1\prec c_{1}\prec b_{2}-1\prec c_{2}$ with $B=\{b_{1},b_{2}\}$ and $C=\{c_{1},c_{2}\}$. Moreover, we have that if $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\neq 0$, then $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{B},\mathsf{O}_{C})\cong K$. Finally, we have $\Sigma\mathsf{O}_{B}=\mathsf{O}_{B-1}$, where $B\pm 1:=\{b\pm 1:b\in B\}$. These descriptions follow, for instance, from specialising the results of [16].
@@ 2.3. Maximal green sequences
To suit our purposes, we use a different presentation of maximal green sequences to the usual one. Our presentation uses the concept of an extriangulated category, for background on which, see [15]. Given a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, we consider the extriangulated structure on $\mathscr{C}_{n}$ given by those triangles $\mathsf{X}\to\mathsf{Y}\to\mathsf{Z}\to\Sigma\mathsf{X}$ for which the morphism $\mathsf{Z}\to\Sigma\mathsf{X}$ factors through $\Sigma\mathsf{T}$. We write $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Z},\mathsf{X})$ for $\operatorname{Ext}$ in this extriangulated structure. This extriangulated structure was first introduced in [17, Section 4.7.1] and is an example of a relative extriangulated structure as studied in [12, Section 3.2].
We say that two basic cluster tilting objects $\mathsf{T}_{1}$ and $\mathsf{T}_{2}$ are related by a mutation if $\mathsf{T}_{1}=\mathsf{E}\oplus\mathsf{X}$ and $\mathsf{T}_{2}=\mathsf{E}\oplus\mathsf{Y}$, where $\mathsf{X}$ and $\mathsf{Y}$ are indecomposable and not isomorphic. We say that the mutation from $\mathsf{T}_{1}$ to $\mathsf{T}_{2}$ is $\mathsf{T}$-green if $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Y},\mathsf{X})\neq 0$. We then say that $(\mathsf{X},\mathsf{Y})$ is the exchange pair of the mutation.
A maximal green sequence of $\mathsf{T}$ is a sequence $\mathcal{G}=(\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l})$ of basic cluster-tilting objects $\mathsf{T}_{i}\in\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{l}=\Sigma\mathsf{T}$, and for all $i\in[l]$, $\mathsf{T}_{i}$ is a $\mathsf{T}$-green mutation of $\mathsf{T}_{i-1}$. Denote by $(\mathsf{X}_{i},\mathsf{Y}_{i})$ the exchange pair of the $\mathsf{T}$-green mutation from $\mathsf{T}_{i-1}$ to $\mathsf{T}_{i}$. We write $\mathrm{Exch}(\mathcal{G}):=\{(\mathsf{X}_{1},\mathsf{Y}_{1}),(\mathsf{X}_{1},\mathsf{Y}_{1}),\dots,(\mathsf{X}_{l},\mathsf{Y}_{l})\}$ for the set of exchange pairs of $\mathcal{G}$. We say that two maximal green sequences $\mathcal{G}$ and $\mathcal{G}^{\prime}$ of $\mathsf{T}$ are equivalent if $\mathrm{Exch}(\mathcal{G})=\mathrm{Exch}(\mathcal{G}^{\prime})$. See [11, Theorem 3.19] for other characterisations of this equivalence relation.
@@
What we call a maximal green sequence of $\mathsf{T}$ corresponds to a maximal green sequence of $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ in the literature. We explain how our definition is equivalent to the definition of maximal green sequence using two-term silting complexes (see [11, Section 2.3.5]).
It follows from [8, Theorem 6.15] that the category $K^{[-1,0]}(\mathrm{proj}\,\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T})$ of two-term complexes of projectives over $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ is equivalent, as an extriangulated category, to the extriangulated ideal quotient of $\mathscr{C}_{n}$ (with the above extriangulated structure) by the morphisms from the shifted projectives to the projectives. It follows from [8, Proposition 5.1] that the relative extriangulated structure used in [8, Theorem 6.15] is the one described above.

### v3

# Oriented matroids from type $\mathbb{A}$ cluster categories

@@ Abstract.
I would like to thank Rudradip Biswas for the opportunity to speak about a preliminary version of this work at the “New Directions in Group Theory and Triangulated Categories” seminar, and Mikhail Gorsky for the reference [FGP+24]. Thank you also to some referees for useful comments. I am currently supported by EPSRC grant EP/W001780/1.
@@ 1. Introduction
From the very beginning, there has been an intimate relationship between cluster algebras and the combinatorics of triangulations. The most fundamental example of this is the $\mathbb{A}_{n}$ cluster algebra [FZ02], whose cluster variables are in bijection with arcs in a convex $(n+3)$-gon, with clusters in bijection with triangulations, and mutation corresponding to flipping an arc in a quadrilateral. This was extended to cluster algebras from triangulations of surfaces in [FST08].
Higher-even-dimensional triangulations were connected with cluster algebras in [OT12], and higher-odd-dimensional ones in [Wil22]. In particular, in [Wil22] it was shown that equivalence classes of maximal green sequences of a linearly oriented type $\mathbb{A}$ quiver were in bijection with triangulations of a three-dimensional cyclic polytope. Maximal green sequences are finite sequences of mutations in a cluster algebra, first introduced into the mathematics literature by Keller [Kel11], and into the physics literature by Cecotti, Cordova, and Vafa [CCV11, Section 4.3]. In fact, these physicists already made a connection between maximal green sequences and three-dimensional triangulations [CCV11, Section 5]. This appearance of maximal green sequences in physics is related to their role in quantum dilogarithm identities [Rei10] and the Kontsevich–Soibelman wall-crossing formula from Donaldson–Thomas theory [KS08].
@@
Given the bijection from [Wil22] between triangulations of three-dimensional cyclic polytopes and equivalence classes of maximal green sequences of linearly-oriented $\mathbb{A}_{n}$, the following natural question then arises. Given an arbitrary orientation of the type $\mathbb{A}$ Dynkin diagram, is there a three-dimensional polytope whose triangulations are in bijection with its maximal green sequences, up to equivalence? Following [Wil22], this bijection should be such that the simplices in the polytope are in bijection with the mutations in the maximal green sequence, so that the number of simplices of the triangulation equals the length of the maximal green sequence.
One can deduce what properties such a three-dimensional polytope should have. An exchange pair in a maximal green sequence corresponds to flipping an arc in a quadrilateral. As in [Wil22], one should lift this quadrilateral to a $3$-simplex such that the green mutation goes from the lower side of the simplex to the upper side. These two ways of lifting a quadrilateral to a $3$-simplex are shown in Figure 1. Choosing whether a $3$-simplex in our hypothetical polytope is like the left-hand simplex in this figure or the right-hand simplex corresponds to choosing an “orientation” for the $3$-simplex. This determines the “oriented matroid” of the hypothetical polytope.
[…]
Figure 1. Two ways to lift a quadrilateral to a $3$-simplex
In order to determine the orientation of each $3$-simplex, we need a criterion which tells us, given an initial cluster, which mutations are green. A convenient criterion is given by [FGP+24] in terms of the cluster category of [BMR+06], where a choice of initial cluster is given by a choice of cluster-tilting object $\mathsf{T}$. In particular, different orientations of $\mathbb{A}_{n}$ can be achieved by different choices of $\mathsf{T}$. Of course, it is not obvious that orienting the $3$-simplices according to which exchange pairs are made green by the cluster-tilting object $\mathsf{T}$ gives a well-defined oriented matroid $\mathcal{M}_{\mathsf{T}}$: certain axioms need to be satisfied. The first main theorem of this paper is that these axioms do in fact hold. We moreover conjecture that $\mathcal{M}_{\mathsf{T}}$ is indeed the oriented matroid of some polytope: that $\mathcal{M}_{\mathsf{T}}$ is “realisable”.
@@
The oriented matroids $\mathcal{M}_{\mathsf{T}}$ are orientations of the uniform matroid: all possible $3$-simplices are bases of the matroid. Orientations of the uniform matroid have been studied recently in the context of their associated “chirotropical Grassmannians”, which were introduced in [CEZ24] in the context of scattering amplitudes, and subsequently studied in [AE24]. While uniform matroids are always realisable [BLVS+99, Section 8.3], uniform oriented matroids are not always realisable [BLVS+99, Proposition 8.3.2].
@@ 2. Background
We start by giving background on oriented matroids, cluster categories, and maximal green sequences.
@@ 2.1. Oriented matroids
A matroid is a structure abstracting the linear dependence relations that hold among a finite set of vectors. An oriented matroid is a matroid with some extra structure — an “orientation” — which also keeps track of the signs of the coefficients in these linear relations. We will follow the standard reference for oriented matroids [BLVS+99].
@@
    $\chi(x_{\pi(1)},x_{\pi(2)},\dots,x_{\pi(r)})=\mathsf{sign}(\pi)\chi(x_{1},x_{2},\dots,x_{r})$
@@
The chirotope $\chi$ is realisable (over $\mathbb{R}$) if there exists a map $\rho\colon E\to\mathbb{R}^{r}$ such that
@@
A basis of $\chi$ is a subset $B=\{b_{1},\dots,b_{r}\}\subseteq E$ such that $\chi(b_{1},\dots,b_{r})\neq 0$. Note that by (2), this is independent of the ordering on $B$. A subset of $E$ is independent if it is contained in a basis, and otherwise it is dependent. A dependent subset of $E$ which is minimal with respect to inclusion is called a circuit.
The set of bases of $\chi$ constitutes the underlying unoriented matroid. If the set of bases is $\binom{E}{r}:=\{B\subseteq E:|B|=r\}$, then the underlying unoriented matroid is the uniform matroid. This will be the case for the matroids that we study in this paper.
The chirotope of an oriented matroid can be used to give the circuits signatures in the following sense. We do not give detail on how the signature of a circuit is derived from the chirotope; the reference is [BLVS+99, Lemma 3.5.7].
Definition 2.2 ([BLVS+99, p.102]).
A signed set $X$ is a set $\underline{X}$ together with a partition $(X^{+},X^{-})$ of $\underline{X}$ into two distinguished subsets. Here $\underline{X}$ is called the support of $X$. We call $(X^{+},X^{-})$ a signature of $X$. Given a set $E$, a signed subset of $E$ is a signed set $X$ with $\underline{X}\subseteq E$.
One can specify axioms for a set of signed subsets of the ground set $E$ to be the set of signed circuits of an oriented matroid [BLVS+99, Definition 3.2.1]. Note that we do not identify an oriented matroid with its chirotope or with its set of signed circuits. We think of oriented matroids as objects in the style of object-oriented programming: an oriented matroid has a chirotope and a set of signed circuits as attributes; it is determined by specifying either of these pieces of data, but it is not identified with either of them. We say that an oriented matroid is realisable (over $\mathbb{R}$) if its chirotope is realisable (over $\mathbb{R}$). The following example of an oriented matroid will occur in a couple of places later in the paper.
Example 2.3.
We denote $m]:=\{1,2,\dots,m\}$. There is a rank-four oriented matroid on $[m]$ defined by the chirotope
$\chi(a,b,c,d)=+1\quad\text{ if }a<b<c<d,$
whose value on quadruples which are not totally ordered is determined by the requirement that $\chi$ is alternating. This is the oriented matroid of a three-dimensional cyclic polytope with $m$ vertices and can be realised using Vandermonde matrices, see [BLVS+99, Section 9.4] for more details. The signed circuits of this oriented matroid are given by $(\{a,c,e\},\{b,d\})$ and $(\{b,d\},\{a,c,e\})$ for $a<b<c<d<e$ [Bre73].
@@
Let $\chi\colon E^{r}\to\{-1,0,+1\}$ be the chirotope of $\mathcal{M}$. If $\mathcal{M}(V)$ also has rank $r$, then its chirotope is $\chi|_{V^{r}}$.
For every oriented matroid $\mathcal{M}$ on a set $E$, there is a dual oriented matroid $\mathcal{M}^{\ast}$ on $E$ [BLVS+99, Proposition 3.4.1]. The (unsigned) circuits of $\mathcal{M}^{\ast}$ are the subsets $Y\subseteq E$ which are inclusion-minimal with respect to the property that $|Y\cap X|\neq 1$ for all circuits $X$ of $\mathcal{M}$. The circuits of $\mathcal{M}^{\ast}$ are known as the cocircuits of $\mathcal{M}$. Lemma 3.13, which we will see later, tells us how the signatures of the cocircuits can be derived from the chirotope.
If $\mathcal{M}$ is realisable, then linear hyperplanes give cocircuits of $\mathcal{M}$: one half of the cocircuit is the subset of $E$ lying on one side of the hyperplane, the other half is the subset lying on the other [BLVS+99, p.116], and the complement of the cocircuit is the subset lying in the hyperplane.
Definition 2.5 ([BLVS+99, p.377]).
@@
Let $\mathcal{M}$ be an acyclic oriented matroid on a set $E$. A hyperplane of $\mathcal{M}$ is a subset $H\subseteq E$ such that $E\setminus H$ is a cocircuit of $\mathcal{M}$. An open halfspace of $\mathcal{M}$ is $Y^{+}$ for $Y$ a cocircuit of $\mathcal{M}$. A facet of $\mathcal{M}$ is a hyperplane $H$ such that $E\setminus H$ is an open halfspace.
@@
    If $\sigma\in\Delta$, then each facet $\tau$ of $\mathcal{M}(\sigma)$ is either contained in a facet of $\mathcal{M}$ or there exists $\sigma^{\prime}\in\Delta\setminus\{\sigma\}$ such that $\tau\subseteq\sigma^{\prime}$.
@@ 2.2. Cluster categories of type $\mathbb{A}$
Cluster categories were introduced in the paper [BMR+06], which we refer to for full background. In type $\mathbb{A}$, these categories were first introduced in [CCS06]. Given an acyclic quiver $Q$, its path algebra over a field $K$ is denoted $KQ$. We write $\mathscr{D}^{b}(KQ)$ for the bounded derived category of right $KQ$-modules which are finite-dimensional over $K$. The cluster category $\mathscr{C}_{Q}$ of $Q$ is the orbit category $\mathscr{D}^{b}(KQ)/\nu\Sigma^{-2}$, where $\nu:=D(KQ)\otimes^{\mathbf{L}}_{KQ}-\colon\mathscr{D}^{b}(KQ)\to\mathscr{D}^{b}(KQ)$ is the Nakayama functor and $\Sigma$ is the suspension of $\mathscr{D}^{b}(KQ)$. Here $D(-)=\operatorname{Hom}_{K}(-,K)$ is the standard duality.
The linearly oriented $\mathbb{A}_{n}$ quiver is $1\leftarrow 2\leftarrow\dots\leftarrow n$ and we denote its cluster category by $\mathscr{C}_{n}$. For technical reasons, we prefer to have the arrows of this quiver going in this direction rather than to the right, which is more usual. (However, $\mathscr{C}_{n}$ is independent of the choice of orientation of the quiver.) The cluster category $\mathscr{C}_{n}$ is a $K$-linear triangulated category with finitely many indecomposable objects, and one can describe its morphisms and suspension functor combinatorially. Indeed, up to isomorphism, the indecomposable objects of $\mathscr{C}_{n}$ may be labelled $\mathsf{O}_{B}$ for $B=\{b_{1},b_{2}\}\in\binom{[n+3]}{2}$ with $b_{1}\neq b_{2}\pm 1$. Here, when we write $b_{2}\pm 1$, we are using modular arithmetic with representatives in $[n+3]$, rather than $\{0,1,\dots,n+2\}$, as is more usual; we will continue do this in this paper. Hence, indecomposable objects in $\mathscr{C}_{n}$ are in bijection with arcs in a convex $(n+3)$-gon.
[…]
@@ 2.3. Maximal green sequences
To suit our purposes, we use a different presentation of maximal green sequences to the usual one. Our presentation uses the concept of an extriangulated category, for background on which, see [NP19]. Given a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, we consider the extriangulated structure on $\mathscr{C}_{n}$ given by those triangles $\mathsf{X}\to\mathsf{Y}\to\mathsf{Z}\to\Sigma\mathsf{X}$ for which the morphism $\mathsf{Z}\to\Sigma\mathsf{X}$ factors through $\Sigma\mathsf{T}$. We write $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Z},\mathsf{X})$ for $\operatorname{Ext}$ in this extriangulated structure. This extriangulated structure was first introduced in [PPPP23, Section 4.7.1] and is an example of a relative extriangulated structure as studied in [HLN21, Section 3.2].
We say that two basic cluster tilting objects $\mathsf{T}_{1}$ and $\mathsf{T}_{2}$ of $\mathscr{C}_{n}$ are related by a mutation if $\mathsf{T}_{1}=\mathsf{E}\oplus\mathsf{X}$ and $\mathsf{T}_{2}=\mathsf{E}\oplus\mathsf{Y}$, where $\mathsf{X}$ and $\mathsf{Y}$ are indecomposable and not isomorphic. We say that the mutation from $\mathsf{T}_{1}$ to $\mathsf{T}_{2}$ is $\mathsf{T}$-green if $\operatorname{Ext}_{\mathsf{T}}^{1}(\mathsf{Y},\mathsf{X})\neq 0$. We then say that $(\mathsf{X},\mathsf{Y})$ is the exchange pair of the mutation.
A maximal green sequence of $\mathsf{T}$ is a sequence $\mathcal{G}=(\mathsf{T}_{0},\mathsf{T}_{1},\dots,\mathsf{T}_{l})$ of basic cluster-tilting objects $\mathsf{T}_{i}\in\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{l}=\Sigma\mathsf{T}$, and for all $i\in[l]$, $\mathsf{T}_{i}$ is a $\mathsf{T}$-green mutation of $\mathsf{T}_{i-1}$. Denote by $(\mathsf{X}_{i},\mathsf{Y}_{i})$ the exchange pair of the $\mathsf{T}$-green mutation from $\mathsf{T}_{i-1}$ to $\mathsf{T}_{i}$. We say that $\mathsf{X}_{i}$ enters $\mathcal{G}$ at $(\mathsf{X}_{i},\mathsf{Y}_{i})$ and $\mathsf{Y}_{i}$ leaves $\mathcal{G}$. We write $\mathrm{Exch}(\mathcal{G}):=\{(\mathsf{X}_{1},\mathsf{Y}_{1}),\dots,(\mathsf{X}_{l},\mathsf{Y}_{l})\}$ for the set of exchange pairs of $\mathcal{G}$. We say that two maximal green sequences $\mathcal{G}$ and $\mathcal{G}^{\prime}$ of $\mathsf{T}$ are equivalent if $\mathrm{Exch}(\mathcal{G})=\mathrm{Exch}(\mathcal{G}^{\prime})$. See [GW23, Theorem 3.19] for other characterisations of this equivalence relation.
@@
What we call a maximal green sequence of $\mathsf{T}$ corresponds to a maximal green sequence of $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ in the literature. We explain how our definition is equivalent to the definition of maximal green sequence using two-term silting complexes (see [GW23, Section 2.3.5]).
It follows from [FGP+24, Theorem 6.15] that the category $K^{[-1,0]}(\mathrm{proj}\,\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T})$ of two-term complexes of projectives over $\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T}$ is equivalent, as an extriangulated category, to the extriangulated ideal quotient of $\mathscr{C}_{n}$ (with the above extriangulated structure) by the morphisms from the shifted projectives to the projectives. It follows from [FGP+24, Proposition 5.1] that the relative extriangulated structure used in [FGP+24, Theorem 6.15] is the one described above. Moreover, this quotient induces a bijection between basic cluster-tilting objects in $\mathscr{C}_{n}$ and two-term silting complexes in $K^{[-1,0]}(\mathrm{proj}\,\operatorname{End}_{\mathscr{C}_{n}}\mathsf{T})$ [FGP+24, Remark 6.20].